OpenStudy (anonymous):

Solve the following system of equations: y = x2 + 3 y = x + 5 Please i just want someone to walk me through the steps of substitution?

OpenStudy (anonymous):

put \[x^2+3=x+5\] and solve the quadratic equation

OpenStudy (anonymous):

I tried that and I know my answer is wrong because it doesn't make any sense

OpenStudy (madhu.mukherjee.946):

x=-1,x=2

OpenStudy (madhu.mukherjee.946):

y=x+5 so put this value of y in the second equation,y=x^2+3 x+5=x^2+3 x^2-x-2=0 now middle term breaking x^2-2x+x-2=0 now take x common from the first part

OpenStudy (madhu.mukherjee.946):

and 1 from the second part so it is ,x(x-2)+1(x-2)=0 (x-2)(x+1)=0 x=2or x=-1

OpenStudy (anonymous):

Thank you so much!!